Write the mathematical expressions for dv  in spherical and cylindrical coordinates.


1
Expert's answer
2021-10-03T13:23:24-0400


First, for the cylindrical coordinates we have

"x=r \cos\theta
\\ y=r\sin\theta
\\ z=z"


Then, we use the jacobian to perform the change of variables "(x,y,z)\to(r,\theta,z)" and for this particular case:


"dV_{cylindrical}=J^* {dr}{d\theta}{dz},
\\ \text{where J* is the jacobian:}

\\ J^*=\text{absolute value of }det \begin{vmatrix}
 x_{r} & x_{\theta} & x_{z} \\
 y_{r} & y_{\theta} & y_{z} \\
 z_{r} & z_{\theta} & z_{z}
\end{vmatrix}
\\ \text{we compute the partial}
\\ \text{derivatives to find:}
\\ J^*= \begin{vmatrix}
 \cos \theta & -r \sin \theta & 0 \\
 \sin \theta & r \cos \theta & 0 \\
 0 & 0 & 1
\end{vmatrix}
\\ J^*= (1)\begin{vmatrix}
 \cos \theta & -r \sin \theta \\
 \sin \theta & r \cos \theta
\end{vmatrix}
\\ \text{ }
\\\implies J^*=r(\cos^2\theta+\sin^2\theta)=r"


This conclusion leads us to: "dV_{cylindrical}=r{dr}{d\theta}{dz}".


Now, for the spherical coordinates, we have


"x=\rho \sin\phi\cos\theta
\\ y=\rho \sin\phi\sin\theta
\\ z=\rho\cos\phi"


and the volume differential (for the new change of variables "(x,y,z)\to(\rho,\phi,\theta)") is defined as


"dV_{spherical}=J {d\rho}{d\phi}{d\theta},
\\ \text{where J is the jacobian:}

\\ J=\text{absolute value of }det \begin{vmatrix}
 x_{\rho} & x_{\phi} & x_{\theta} \\
 y_{\rho} & y_{\phi} & y_{\theta} \\
 z_{\rho} & z_{\phi} & z_{\theta}
\end{vmatrix}
\\ \text{we compute the partial}
\\ \text{derivatives to find:}
\\ J= \begin{vmatrix}
 \sin \phi \cos \theta & \rho \cos \phi \cos \theta & - \rho \sin \phi \sin \theta \\
 \sin \phi \sin \theta & \rho \cos \phi \sin \theta & \rho \sin \phi \cos \theta \\
 \cos \phi & - \rho \sin \phi & 0
\end{vmatrix}
\\ \text{ }
\\ \implies J= \cos\phi \begin{vmatrix}
 \rho \cos \phi \cos \theta & - \rho \sin \phi \sin \theta \\
 \rho \cos \phi \sin \theta &\rho \sin \phi \cos \theta
\end{vmatrix}+\rho \sin\phi \begin{vmatrix}
 \sin \phi \cos \theta & - \rho \sin \phi \sin \theta \\
 \sin \phi \sin \theta & \rho \sin \phi \cos \theta 
\end{vmatrix}"


Then we continue to reduce and solve the determinants to find J:


"J=\cos \phi (\rho^2\sin\phi\cos\phi\cos^2\theta+\rho^2\sin\phi\cos\phi\sin^2\theta)+...
\\ ...+(\rho\sin\theta)( \rho\sin^2\phi\cos^2\theta+\rho\sin^2\phi\sin^2\theta )
\\ \text{once we reduce and use trigonometric identities},
\\ \text{ we find } J=\rho^2 \sin \phi"


Thus we finally have: "dV_{spherical}=\rho^2\sin\theta {d\rho}{d\phi}{d\theta}"



Reference:

  • Department of Mathematics. (1996). Triple Integrals in Cylindrical and Spherical Coordinates. Triple integrals in cylindrical and spherical coordinates. Retrieved October 3, 2021, from http://sites.science.oregonstate.edu/math/home/programs/undergrad/CalculusQuestStudyGuides/vcalc/255cs/255cs.html